OpenStudy (anonymous):

John ran 8 miles in 1 hour. At that rate, how many miles will he run in 2 hours? A. 16 mi B. 10 mi C. 4 mi D. 2 mi

hero (hero):

\[\frac{8 \space\text{miles}}{x \space\text{miles}} = \frac{1 \space\text{hour}}{2 \space\text{hour}}\] \[\frac{8}{x} = \frac{1}{2}\]

OpenStudy (anonymous):

^^ 1x=16 1x/1=x, and 16/1=16 x=16

OpenStudy (anonymous):

10

OpenStudy (anonymous):

What? The answer is 16

OpenStudy (anonymous):

ok

hero (hero):

If \[\frac{a}{b} = \frac{c}{d}\] then \[(a)(d) = (b)(c)\]